# FLOPS: Forward Learning with OPtimal Sampling

### v1

@@ Abstract
Given the limitations of backpropagation, perturbation-based gradient computation methods have recently gained focus for learning with only forward passes, also referred to as queries. Conventional forward learning consumes enormous queries on each data point for accurate gradient estimation through Monte Carlo sampling, which hinders the scalability of those algorithms. However, not all data points deserve equal queries for gradient estimation. In this paper, we study the problem of improving the forward learning efficiency from a novel perspective: how to reduce the gradient estimation variance with minimum cost? For this, we propose to allocate the optimal number of queries over each data in one batch during training to achieve a good balance between estimation accuracy and computational efficiency. Specifically, with a simplified proxy objective and a reparameterization technique, we derive a novel plug-and-play query allocator with minimal parameters. Theoretical results are carried out to verify its optimality. We conduct extensive experiments for fine-tuning Vision Transformers on various datasets and further deploy the allocator to two black-box applications: prompt tuning and multimodal alignment for foundation models. All findings demonstrate that our proposed allocator significantly enhances the scalability of forward-learning algorithms, paving the way for real-world applications.
@@ 1 Introduction
Ever since the success of backpropagation (BP) (Rumelhart et al., 1986), researchers have sought alternate biologically plausible methods as the brain doesn’t perform backward computation to learn high-level representation (Ma et al., 2020; Nøkland, 2016; Jacot et al., 2018). Moreover, existing practical scenarios urge for scalable forward learning methods, when the black-box nature renders taking the derivative infeasible or difficult. For example, as the model parameters increase, machine learning (ML) systems usually integrate with third-party black-box APIs (Achiam et al., 2023).
@@
However, it is not always helpful for the forward learning algorithms to increase the number of queries. Especially for the large model training on a large dataset, it usually requires an extensive number of queries, which poses great challenges to the memory and computation cost. For example, The forward-forward (FF) (Hinton, 2022) algorithm and the feedback alignment (FA) (Nøkland, 2016) are only capable of training multilayer perceptrons on MNIST (LeCun, 1998). To scale up the forward algorithm to large model training on large datasets, DeepZero (Chen et al., 2023) and Mezo (Malladi et al., 2023) explore the use of SPSA (Spall, 1992) in training the convolution network from scratch and fine-tune the language model, respectively. But with such “large” scale optimization problem, DeepZero requires $192$ queries to train a pruned ResNet-20 with only $12K$ parameters. Mezo needs meticulously selected prompts to ensure performance, and its gradient estimator is too noisy due to insufficient query numbers, thus suffering from highly unstable performance.
@@
  3. 3.
    Extensive experiments are conducted on fine-tuning Vision Transformer (Dosovitskiy, 2020) on various challenging datasets including the ImageNet-1K Deng et al. (2009), achieving the SOTA performance in the realm of forward learning.
@@
    We leverage the proposed method to solve two real-world black-box problems, including black-box prompt tuning and alignment between foundation models, showing its scalability and efficiency improvement versus a series of optimization methods, including the BP.
@@ 2 Related Work
Backpropagation-free learning. Without BP, people have designed algorithms drawing inspiration from biological social behavior and evolution theory Wierstra et al. (2014). Swarm intelligent algorithms, for example, particle swarm optimization (Vaz & Vicente, 2009), optimize a black-box problem by iteratively tracking the best candidate in the population. Evolutionary algorithms, for example, natural evolutionary strategies, inject random noise into the model and optimize along the optimal direction along the noise (Salimans et al., 2017). Both swarm intelligence algorithms and evolution algorithms lack sufficient theoretical analysis and the performance would drop as the parameters increase. Forward gradient (Baydin et al., 2022; Silver et al., 2021), applying the chain rule from a different direction than BP, utilizes forward-mode automatic differentiation (AD) to train the ML model. Forward-mode AD needs specific AD software and full access to the model structure. Therefore, training the deep model under black-box scenarios is not feasible by the forward gradient (Ren et al., 2022).
Perturbation-based forward learning. Besides the heuristic algorithms, researchers have devised surrogate methods to approximate the first-order gradients, such as SPSA. The SPSA injects noise with opposite signs into the model parameters to form a finite-difference style estimator. Recently, Mezo (Malladi et al., 2023) applied SPSA to fine-tune the language model. DeepZero (Chen et al., 2023) integrates SPSA with sparsity and trains the network from scratch with many queries. However, their methods still have some limitations and their query budgets are equally allocated, neglecting the data difference. Jiang et al. (2023) extends the likelihood ratio estimator to train a wide range of neural networks, i.e. CNN (He et al., 2016), RNN (Dey & Salem, 2017), GNN (Scarselli et al., 2008), SNN (Ghosh-Dastidar & Adeli, 2009), whereas they don’t verify their method on modern architecture, i.e., Transformers (Vaswani, 2017). There is still a huge research gap for improving the scalability of forward learning.
@@
3 Preliminary on Forward Learning
@@
Given a generic neural network with no assumption on its architecture, the input is denoted as $x\in\mathbb{R}^{d_{x}}$, and the output is given by $y=\phi_{2}(\varphi(\phi_{1}(x);\theta))\in\mathbb{R}^{d_{y}}$, where $\phi_{1}(\cdot)$ and $\phi_{2}(\cdot)$ are black-box parts, and $\varphi(\cdot;\theta)$ is the trainable structure parameterized by $\theta\in\Theta\subset\mathbb{R}^{d_{\theta}}$. Training the neural network is to solve such an optimization: $\min_{\theta\in\Theta}\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\mathcal{L}(y)$, where $\mathcal{X}$ is the dataset for training, and $\mathcal{L}(\cdot)$ is an evaluation procedure yielding the loss feedback. Due to the growing scale of parameters in modern neural networks and the dataset, stochastic gradient descent has essentially become the only feasible solution method. The primary challenge remains in computing the gradients for each batch efficiently.
@@
In addition to approaches relying on computation graphs like BP, forward-learning algorithms estimate gradients by perturbing the neural activity and observing the correlation between the injected perturbation and the final loss value, thereby solving the stochastic version of the problem: $\min_{\theta\in\Theta}\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\mathbb{E}[\mathcal{L}(y)]$. For instance, Jiang et al. (2023) perturb the intermediate result to perform LR estimation, i.e., $y=\phi_{2}(\varphi(\phi_{1}(x);\theta)+z)$, where $z$ is the injected noise with a density $f(\cdot)$. Denote the Jacobian matrix of $a\in\mathbb{R}^{d_{a}}$ with respect to $x\in\mathbb{R}^{d_{b}}$ as $D_{b}a\in\mathbb{R}^{d_{a}\times d_{b}}$. The loss gradient can be estimated by
$\nabla_{\theta}\mathbb{E}[\mathcal{L}(y)]=\mathbb{E}[-\mathcal{L}(y)D_{\theta}^{\top}\varphi(\phi_{1}(x);\theta)\nabla_{\theta}\ln f(z)].$ (1)
@@
The SPSA family injects noise to the parameters (Spall, 1992; Salimans et al., 2017), i.e., $y=\phi_{2}(\varphi(\phi_{1}(x);\theta+z))$. For continuous noise distributions, the gradient estimation can be derived as a special case of equality (1) as follows:
$\nabla_{\theta}\mathbb{E}[\mathcal{L}(y)]=\mathbb{E}[-\mathcal{L}(y)\nabla_{\theta}\ln f(z)].$ (2)
@@
It can also be interpreted as a method for randomized finite differences in high-dimensional space if we use Gaussian noise $z\sim\mathcal{N}(0,\sigma^{2}I_{d})$ and apply an antithetic variable trick for variance reduction to equality (2), i.e.,
$\nabla_{\theta}\mathbb{E}[\mathcal{L}(y)]=\mathbb{E}\big{[}\frac{z}{2\sigma}(\mathcal{L}(\phi_{2}(\varphi(\phi_{1}(x);\theta+z))-\mathcal{L}(\phi_{2}(\varphi(\phi_{1}(x);\theta-z)))\big{]}.$
@@
Both equations (1) and (2) use Monte-Carlo simulation with $A>0$ queries to compute the loss gradient, and the corresponding estimates take the common form $\nabla_{\theta}\hat{\mathcal{L}}(y)=\frac{1}{A}\sum_{i=1}^{A}G(x,\theta,z_{i})$, where $z_{i}$ is sampled independently from $f(\cdot)$.
@@
Given the limited computational resources, the parameter is updated iteratively with randomly sampled mini-batches of data denoted as $\{x_{1},\cdots,x_{B}\}$, where $B$ is the batch size. For readability, we abbreviate $\nabla_{\theta}\hat{\mathcal{L}}(y_{j})=\frac{1}{A_{j}}\sum_{i=1}^{A_{j}}G(x_{j},\theta,z_{i,j})$ as $\nabla_{\theta}\hat{\mathcal{L}}_{j}(\theta)$ and denote the batch-wise gradient estimates as $\nabla_{\theta}\hat{\mathcal{L}}(\theta_{t})$. The updated recursion can be written as
$\theta_{t+1}=\theta_{t}-\eta_{t}\nabla_{\theta}\hat{\mathcal{L}}(\theta_{t})=\theta_{t}-\frac{\eta_{t}}{B}\sum_{j=1}^{B}\nabla_{\theta}\hat{\mathcal{L}}_{j}(\theta_{t}),$ (3)
@@
By default, the query amount $A_{j}$ allocated to estimate the gradient associated with each data point is equally allocated. However, the difficulty of gradient estimation and the model performance differ over data points. To conserve computational resources while maintaining the quality of gradient estimation, the number of queries should be properly allocated across data within the mini-batch. Utilizing the likelihood ratio technique, we propose a universal accelerator for all perturbation-based training methods from a data-aware perspective.
4 Method
4.1 Optimal allocation via likelihood ratio method
@@
Bernoulli Allocator (BA) is another lightweight alternative, similar to the idea in Qin et al. (2023). In an equal allocation scenario, each data in the batch receives $\overline{A}$ queries. The BA introduces a probabilistic mechanism where, with probability $p$, the number of queries assigned to a data point is reduced to $\overline{A}/2$; otherwise, it remains at $\overline{A}$. The probability $p$ is the only parameter of the allocator.
@@
Optimization of the Query Allocator. We employ a gradient-based method to optimize the allocator parameters $\bm{\lambda}$, introducing an additional optimization step before each iteration of neural network training. However, by formulating an equivalent optimization objective and utilizing reparameterization techniques, we ensure that this auxiliary optimization remains computationally lightweight and incurs minimal overhead.
@@
The LR gradient estimator
@@
where $\bm{A}^{(k)}=(A_{1}^{(k)},\cdots,A_{B}^{(k)})$ are i.i.d samples from $P(\cdot|\bm{\lambda})$. At each step $t$ of neural network training, we employ the above gradient estimator to search for the optimal allocator parameter $\bm{\lambda}_{t}^{\ast}\triangleq\arg\max_{\bm{\lambda}\in\Lambda}J(\bm{\lambda})$. Let $LB^{\ast}_{1:T}$ and $LB_{1:T}^{\mathrm{equal}}$ represent the lower bound of the cumulative performance improvement of the neural network after $T$ steps of training when using the optimal allocator $\bm{A}^{\ast}\sim P(\cdot|\lambda^{\ast}_{t})$ and using an equal allocator $\bm{A}^{\mathrm{equal}}$, respectively. Then the difference between these two allocation strategies is lower bound by the following expression.
@@
Reparameterization of the Gaussian allocator. The original parameter dimension of the Gaussian allocator is $B+\frac{B(B+1)}{2}$, which might be problematic due to its high dimensionality. To reduce the dimensionality of the allocator’s optimization problem and to incorporate more relevant features, we adopt a reparameterization approach, i.e., imposing a specific structure to the parameter $\bm{\lambda}$.

### v3

@@ Abstract
Given the limitations of backpropagation, perturbation-based gradient computation methods have recently gained focus for learning with only forward passes, also referred to as queries. Conventional forward learning consumes enormous queries on each data point for accurate gradient estimation through Monte Carlo sampling, which hinders the scalability of those algorithms. However, not all data points deserve equal queries for gradient estimation. In this paper, we study the problem of improving the forward learning efficiency from a novel perspective: how to reduce the gradient estimation variance with minimum cost? For this, we allocate the optimal number of queries within a set budget during training to balance estimation accuracy and computational efficiency. Specifically, with a simplified proxy objective and a reparameterization technique, we derive a novel plug-and-play query allocator with minimal parameters. Theoretical results are carried out to verify its optimality. We conduct extensive experiments for fine-tuning Vision Transformers on various datasets and further deploy the allocator to two black-box applications: prompt tuning and multimodal alignment for foundation models. All findings demonstrate that our proposed allocator significantly enhances the scalability of forward-learning algorithms, paving the way for real-world applications. The implementation is available at https://github.com/RTkenny/FLOPS-Forward-Learning-with-OPtimal-Sampling
@@ 1 Introduction
Ever since the success of backpropagation (BP) (Rumelhart et al., 1986), researchers have sought alternate methods to bypass the iterative computation of the backward pass for the sake of efficiency, interpretability, and biological plausibility (Ma et al., 2020; Nøkland, 2016; Jacot et al., 2018). Moreover, existing practical scenarios urge for scalable forward learning methods, when the black-box nature renders taking the derivative infeasible or difficult. For example, as the model parameters increase, machine learning (ML) systems usually integrate with third-party black-box APIs (Achiam et al., 2023).
@@
However, it is not always helpful for the forward learning algorithms to increase the number of queries. Especially for the large model training on a large dataset, it usually requires an extensive number of queries, which poses great challenges to the memory and computation cost. For example, The forward-forward (FF) (Hinton, 2022) algorithm and the feedback alignment (FA) (Nøkland, 2016) are only capable of training multilayer perceptrons on MNIST (LeCun, 1998) or CIFAR (Krizhevsky et al., 2009). To scale up the forward algorithm to large model training on large datasets, DeepZero (Chen et al., 2023) and Mezo (Malladi et al., 2023) explore the use of simultaneous perturbation stochastic approximation (SPSA) (Spall, 1992) in training the convolution network from scratch and fine-tune the language model, respectively. But with such “large” scale optimization problem, DeepZero requires $192$ queries to train a pruned ResNet-20 with only $12K$ parameters. Mezo needs meticulously selected prompts to ensure performance, and its gradient estimator is too noisy due to insufficient query numbers, thus suffering from highly unstable performance.
@@
    We unify different BP-free methods under the framework of perturbation-based optimization. We are the first to formulate the query allocation problem for forward learning in the literature.
  2. 2.
@@
    We conduct extensive experiments to fine-tune Vision Transformer (Dosovitskiy, 2020) on challenging datasets, including ImageNet-1K Deng et al. (2009), achieving state-of-the-art performance in forward learning. Furthermore, we apply the proposed method to two real-world black-box problems: black-box prompt tuning and alignment between foundation models. Our approach demonstrates scalability and efficiency improvements over all baselines, including backpropagation (BP).
@@ 2 Related Work
Backpropagation-free learning. People have explored various algorithms without backpropagating the error (Wierstra et al., 2014; Frenkel et al., 2021; Journé et al., 2022; Kao & Hariharan, 2024). Frenkel et al. (2021) eliminates the feedback pathway by leveraging fixed random projections of target labels to hidden layers. Kao & Hariharan (2024) employs a dual-network structure to propagate input and target signals in anti-parallel directions. Forward gradient (Baydin et al., 2022; Silver et al., 2021), applying the chain rule from a different direction than BP, utilizes forward-mode automatic differentiation (AD) to train the ML model. Forward-mode AD needs specific AD software and full access to the model structure. Therefore, training the deep model under black-box scenarios is not feasible by the forward gradient (Ren et al., 2022).
Perturbation-based forward learning. Built on the theory of stochastic optimization, researchers have devised surrogate methods to approximate the first-order gradients, such as SPSA (Spall, 1992). The SPSA injects noise with opposite signs into the model parameters to form a finite-difference style estimator. Recently, Mezo (Malladi et al., 2023) applied SPSA to fine-tune the language model. DeepZero (Chen et al., 2023) integrates SPSA with sparsity and trains the network from scratch with many queries. However, their methods still have some limitations and their query budgets are equally allocated, neglecting the data difference. Jiang et al. (2023) extends the likelihood ratio (LR) estimator to train a wide range of neural networks, whereas LR has not been verified on modern-scale architectures, such as Transformers (Vaswani, 2017), leaving a significant gap for further research to improve the scalability of forward learning.
@@
Advanced technique for data sampling and reshuffling. In the literature on stochastic optimization, apart from uniform sampling, people have designed various methods to sample data points from the dataset by importance sampling (Zhao & Zhang, 2015), reshuffling (Lu et al., 2022), etc. Their method can reduce the variance of the stochastic gradient. Their methods share some similarities with our OPtimal Sampling. However, the key distinction between our work and the existing methods lies in the probability spaces considered. While the existing methods focus on reducing variance in the probability space associated with ”sampling data points from the dataset”, our approach focuses on perturbation-based forward learning, where the variance lies in the probability space of the injected noise for gradient estimation.
3 A Unified Perspective of Perturbation-based Forward Learning
@@
Given a generic neural network with no assumption on its architecture, the input is denoted as $x\in\mathbb{R}^{d_{x}}$, and the output is given by $y=\phi_{2}(\varphi(\phi_{1}(x);\theta))\in\mathbb{R}^{d_{y}}$, where $\phi_{1}(\cdot)$ and $\phi_{2}(\cdot)$ are black-box parts, and $\varphi(\cdot;\theta)$ is the targeted module with trainable parameters $\theta\in\Theta\subset\mathbb{R}^{d_{\theta}}$. Conducting backpropagation through $\phi_{1}(\cdot)$ and $\phi_{2}(\cdot)$ is infeasible or difficult. Training the neural network is to solve such an optimization problem: $\min_{\theta\in\Theta}\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\mathcal{L}(y)$, where $\mathcal{X}$ is the dataset for training, and $\mathcal{L}(\cdot)$ is an evaluation procedure yielding the loss feedback. The growing scale of parameters in modern neural networks and the dataset makes gradient-based methods the only feasible solution. Since forward learning do not rely on the chain rules to compute the true gradient, the primary challenge remains in estimating the gradient accurately and efficiently.
@@
In addition to approaches requiring computation graphs like BP, forward-learning algorithms estimate gradients by perturbing the neural activity (intermediate output or parameter of the targeted module $\varphi(\cdot;\theta)$) and observing the impact of the injected perturbation on the final loss value, thereby solving the stochastic version of the problem: $\min_{\theta\in\Theta}\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\mathbb{E}_{z\sim f(\cdot)}[\mathcal{L}(y)]$, where $z$ is the injected noise with a density $f(\cdot)$.
Forward learning algorithms can be divided into two categories: the LR family and the SPSA/ES family. The LR family perturb the intermediate output of the targeted module to perform estimation Jiang et al. (2023), i.e., $y=\phi_{2}(\varphi(\phi_{1}(x);\theta)+z)$. The gradient can be estimated by
$\nabla_{\theta}\mathbb{E}[\mathcal{L}(y)]=\mathbb{E}_{z\sim f(\cdot)}[G_{LR}(x,\theta,z)]\triangleq\mathbb{E}_{z\sim f(\cdot)}[-\mathcal{L}(y)D_{\theta}^{\top}\varphi(\phi_{1}(x);\theta)\nabla_{z}\ln f(z)],$ (1)
where $D_{b}a\in\mathbb{R}^{d_{a}\times d_{b}}$ denote the Jacobian matrix of $a\in\mathbb{R}^{d_{a}}$ with respect to $b\in\mathbb{R}^{d_{b}}$.
@@
The SPSA/ES family injects noise to the parameters (Spall, 1992; Salimans et al., 2017), i.e., $y=\phi_{2}(\varphi(\phi_{1}(x);\theta+z))$. For continuous noise distributions, the gradient estimation can be derived as a special case of equality (1) as follows:
$\nabla_{\theta}\mathbb{E}[\mathcal{L}(y)]=\mathbb{E}_{z\sim f(\cdot)}[G_{ES}(x,\theta,z)]\triangleq\mathbb{E}_{z\sim f(\cdot)}[-\mathcal{L}(y)\nabla_{z}\ln f(z)],$ (2)
@@
which takes the same form termed as the evolutionary strategies (ES) estimation (Salimans et al., 2017). The SPSA is a special case of ES which uses Gaussian noise $z\sim\mathcal{N}(0,\sigma^{2}I_{d})$ and applies an antithetic variable trick for variance reduction to equality (2). The SPSA can be interpreted as randomized finite differences in high-dimensional space. It takes the form
$\nabla_{\theta}\mathbb{E}[\mathcal{L}(y)]$ $=\mathbb{E}_{z\sim\mathcal{N}(\cdot)}[G_{SPSA}(x,\theta,z)]$ (3)
$=\mathbb{E}_{z\sim\mathcal{N}(\cdot)}\big{[}\frac{z}{2\sigma}(\mathcal{L}(\phi_{2}(\varphi(\phi_{1}(x);\theta+z))-\mathcal{L}(\phi_{2}(\varphi(\phi_{1}(x);\theta-z)))\big{]}.$
@@
To estimate the expectation of gradient according to equation (1, 2 and 3), both SPSA/ES and LR family use Monte Carlo sampling. With $A>0$ queries of Monte-Carlo sampling to estimate the gradient on data point $x$, the corresponding estimations take the sample-average form $\nabla_{\theta}\hat{\mathcal{L}}(y)=\frac{1}{A}\sum_{i=1}^{A}G(x,\theta,z_{i})$, where $z_{i}$ is sampled independently from $f(\cdot)$ and $G(x,\theta,z_{i})$ is the $i$ th sample of the corresponding gradient estimator ($G(x,\theta,z_{i})$ can be $G_{LR}(x,\theta,z_{i})$, $G_{ES}(x,\theta,z_{i})$ or $G_{SPSA}(x,\theta,z_{i})$).
@@
The parameter is updated iteratively with a mini-batch of data points randomly drawn from $\mathcal{X}$, denoted as $\{x_{1},\cdots,x_{B}\}$, where $B$ is the batch size. We denote the gradient estimation on data $x_{j}$ as $\nabla_{\theta}\hat{\mathcal{L}_{j}}(\theta)=\frac{1}{A_{j}}\sum_{i=1}^{A_{j}}G(x_{j},\theta,z_{i,j})$, where $A_{j}$ denotes the allocated query size, and the batch-wise gradient estimation is $\nabla_{\theta}\hat{\mathcal{L}}(\theta_{t})=\frac{1}{B}\sum_{j=1}^{B}\nabla_{\theta}\hat{\mathcal{L}}_{j}(\theta_{t})$. The updating recursion can be written as
$\theta_{t+1}=\theta_{t}-\eta_{t}\nabla_{\theta}\hat{\mathcal{L}}(\theta_{t})=\theta_{t}-\eta_{t}\frac{1}{B}\sum_{j=1}^{B}\nabla_{\theta}\hat{\mathcal{L}}_{j}(\theta_{t})=\theta_{t}-\eta_{t}\frac{1}{B}\sum_{j=1}^{B}\frac{1}{A_{j}}\sum_{i=1}^{A_{j}}G(x_{j},\theta,z_{i,j}),$ (4)
@@
By default, the query amount $A_{j}$ allocated to estimate the gradient associated with each data point $x_{j}$ is equally allocated. However, the difficulty of gradient estimation and the model performance differ over data points. To conserve computational resources while maintaining the quality of gradient estimation, the number of queries should be properly allocated across data within the mini-batch. Utilizing the likelihood ratio technique, we propose a universal accelerator for all perturbation-based training methods from a data-aware perspective.
4 OPtimal Sampling via likelihood ratio method
4.1 Gaussian Allocator and Bernoulli Allocater
@@
Bernoulli Allocator (BA) is another lightweight alternative, similar to the idea in Qin et al. (2023). In an equal allocation scenario, each data in the batch receives $\overline{A}$ queries. The BA introduces a probabilistic mechanism where, with probability $p$, the number of queries assigned to a data point is reduced to $\overline{A}/2$; otherwise, it remains at $\overline{A}$. The probability $p$ is the only parameter of the allocator. The BA is in the form of a conditional probability distribution
$P(A_{j}=\overline{A}/2|\Phi=\mathcal{L}_{j})=\left\{\begin{array}[]{ll}p,&\mathcal{L}_{j}<\overline{\mathcal{L}}\\ 0,&\mathcal{L}_{j}\geq\overline{\mathcal{L}}\end{array}\right.,$ (6)
where $\mathcal{L}_{j}$ is the clean loss on data $x_{j}$, and $\overline{\mathcal{L}}$ is the mean of loss value in the batch(we use the loss value without injected noise into the network).
@@
4.2 Optimization of the Query Allocator.
We employ a gradient-based method to optimize the allocator parameters $\bm{\lambda}$, introducing an additional optimization step before each iteration of neural network training. However, by formulating an equivalent optimization objective and utilizing reparameterization techniques, we ensure that this auxiliary optimization remains computationally lightweight and incurs minimal overhead.
@@
The LR gradient estimator for the allocator takes the sample average form
@@
where $\bm{A}^{(k)}=(A_{1}^{(k)},\cdots,A_{B}^{(k)})$ are i.i.d samples from $P(\cdot|\bm{\lambda})$ and $K$ is the number of sampling from $P(\cdot|\bm{\lambda})$. At each step $t$ of neural network training, we employ the above gradient estimator to search for the optimal allocator parameter $\bm{\lambda}_{t}^{\ast}\triangleq\arg\max_{\bm{\lambda}\in\Lambda}J(\bm{\lambda})$. Let $LB^{\ast}_{1:T}$ and $LB_{1:T}^{\mathrm{equal}}$ represent the lower bound of the cumulative performance improvement of the neural network after $T$ steps of training when using the optimal allocator $\bm{A}^{\ast}\sim P(\cdot|\lambda^{\ast}_{t})$ and using an equal allocator $\bm{A}^{\mathrm{equal}}$, respectively. Then the difference between these two allocation strategies is lower bound by the following expression.
@@
4.3 Reparameterization of the Gaussian allocator.
The original parameter dimension of the Gaussian allocator is $B+\frac{B(B+1)}{2}$, which might be problematic due to its high dimensionality. To reduce the dimensionality of the allocator’s optimization problem and to incorporate more relevant features, we adopt a reparameterization approach, i.e., imposing a specific structure to the parameter $\bm{\lambda}$.